Taxation Question, need a process of solution
20) Bart has the following income and expenses for his Schedule C 1099-MISCs received: Jacks Builders Smiths Fine Homes Murphys Builders Total Revenue Wages Payroll taxes Payroll tax penalties License fee Supplies Cell phone Meals Internet service S 42,000 35,750 24,730 S 102,480 (46,000) (3,519) (300) (200) (5,300 ) (588) (3,000) (780 ) What is Barts self-employment income?
Answer #1

Total revenue of Bart = $102480

Wages payment to staff = 46000

Payroll taxes = $3519

and payroll tax penalties= $300

Here payroll taxes are deducted from wages of staff but payroll tax penalties is to be paid by employer Bart.

Therefore, net wage payment = 46000-3519= $42481

Miscellaneous expenses of Bart = license fee +Supplies+ Cellphone + Meals +Internet Service + Payroll tax penalties

                                              = 200+5300+588+3000+780+300= $10168

Therefore, Bart's self-employment income = Total revenue- net wage payment -Miscellaneous expenses

                                                           = 102480-42481-10168= $49831